# Demystifying LLM-as-a-Judge: Analytically Tractable Model for Inference-Time Scaling

### v1

@@ 1 Introduction
Scaling training compute via larger models and more data drives dramatic gains (Hestness et al., 2017; Kaplan et al., 2020; Hoffmann et al., 2022). In parallel, across tasks, allowing models to ‘think longer’ at inference by sampling multiple candidates, re-ranking with a reward, or aggregating votes consistently improves precision and reliability (Wang et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Wu et al., 2024; Snell et al., 2025). Best-of-$k$ (choose the highest-reward sample based on the feedback of a judge LLM) and majority voting (choose the consensus among the generated answers) have become standard inference-time tools (Brown et al., 2024; Schaeffer et al., 2025a; Chen et al., 2024a; Huang et al., 2025a) along with self-verification (Saunders et al., 2022; Weng et al., 2023).
@@
This simple model recapitulates existing findings on large language model inference and yields novel predictions that we evaluate in this paper. It reproduces the empirical observation that unbounded increases in the number of inference-time samples do not confer additional benefit (Snell et al., 2025). Furthermore, it predicts the existence of an optimal temperature for the reward process. Further, strong rewards shift optimal temperature to a lower value. We empirically examine and confirm this prediction using Meta-Llama-3-8B-Instruct with Mistral-7B-Instruct-v0.3 serving as the judging model.
@@
  • In order to derive analytic insights about the inference time optimization, we derive a series expansion for $\delta$ around large $T$, making explicit its dependence on $n$, $k$, and the alignment between $\mathbf{w}_{R}$ and $\mathbf{w}_{T}$. The analytical expansion shows a sharp dependence on reward quality: when $\mathbf{w}_{R}$ is sufficiently close to $\mathbf{w}_{T}$, i.e., small $\|\mathbf{w}_{R}-\mathbf{w}_{T}\|/\|\mathbf{w}_{T}\|$, increasing $k$ monotonically decreases $\delta$, and the reward $\mathbf{w}_{R}$ that optimizes inference-time selection generally differs from the data-generating teacher $\mathbf{w}_{T}$. In contrast, when $\mathbf{w}_{R}$ is poorly aligned, $\delta$ is non-monotone in $k$, yielding an optimal finite $k$ (see Figure 1), echoing phenomena observed empirically in large language models (Snell et al., 2025). Furthermore we show that at fixed $k$, there exists an optimal temperature $T$ for the rewarding process. Similar yet distinct observation has been made in large language models (Du et al., 2025) for sampling temperature of the model itself. We emphasize that our result is about the sampling temperature of the rewarding process - different from the sampling temperature of the language model.
@@
  • To test our predictions in realistic settings, we have generated inference time samples from Meta-Llama-3-8B-Instruct on openai/gsm8k validation dataset (prompt included $8$ chain of thought demonstrations). For each question and response pair we used Mistral-7B-Instruct-v0.3 to generate a reward score. Finally, we observed the existence of an optimal $k,T$ similar to the theoretical predictions above (see Figure  7).
@@ 1.1 Related works
A growing body of work investigates how to allocate and exploit inference-time compute to improve predictive performance, with empirical gains reported across tasks and domains based on majority voting  (Chen et al., 2024a; Snell et al., 2025; Setlur et al., 2025; Arora and Zanette, 2025; Wu et al., 2024; Liu et al., 2025; Du et al., 2025; Huang et al., 2025a) or a best-of-$k$ strategy  (Wang et al., 2023; Yao et al., 2023a; Brown et al., 2024; Levi, 2024; Schaeffer et al., 2025b; Huang et al., 2025b; Chen et al., 2024b; Du et al., 2025; Chen et al., 2025). These procedures are often paired with reasoning-oriented prompting and structured search that expand the candidate set before selection (Wei et al., 2022; Yao et al., 2023b). The work of  Chen et al. (2024a) presented a theoretical model for majority voting in the premise of classification problems. More close to our work is the theoretical model of Levi (2024) on best-of-$k$ strategy. For a given trained model, both these works explain some of the empirically observed patterns at inference. We study similar questions for the regression model and discuss trade-off between training and inference time compute taking into account the quality of the reward and the sampling process.
@@ 2.1 Training method - prior and posterior distribution
$r(y,\mathbf{x})=-\left(y-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}$ (6)
@@
$\delta=\mathbb{E}_{\mathbf{x}}(\delta(\mathbf{x})),\quad\delta(\mathbf{x})=\mathbb{E}_{y_{1},\dots,y_{k}}\left[\frac{\sum_{i=1}^{k}\big(y_{i}-\mu_{T}(\mathbf{x})\big)^{2}\,e^{-\big(y_{i}-\mu_{R}(\mathbf{x})\big)^{2}/T}}{\sum_{j=1}^{k}e^{-\big(y_{j}-\mu_{R}(\mathbf{x})\big)^{2}/T}}\right].$ (7)
@@ 3.1 Asymptotic Behavior of the Generalization Error via Deterministic Equivalents
$\delta(\mathbf{x})=\mathbb{E}_{y_{i}\sim\mathcal{N}(m(\mathbf{x}),s(\mathbf{x})^{2}),i=1,2,\dots,k}\left[\frac{\sum_{i=1}^{k}\big(y_{i}-\mu_{T}(\mathbf{x})\big)^{2}\,e^{-\big(y_{i}-\mu_{R}(\mathbf{x})\big)^{2}/T}}{\sum_{j=1}^{k}e^{-\big(y_{j}-\mu_{R}(\mathbf{x})\big)^{2}/T}}\right],$ (8)
to the leading order in $\sigma$. Here the posterior predictive has a mean $m(\mathbf{x})$ and variance $\Sigma(\mathbf{x})=s(\mathbf{x})^{2}$ as follows
@@
$\hat{R}=R\Big(1-\alpha\,m_{\bm{\Sigma}}(R)\Big)=\frac{\sigma^{2}}{\gamma^{2}}\alpha,\qquad m_{\bm{\Sigma}}(R):=\frac{1}{d}\text{Tr}\!\big[\bm{\Sigma}(\bm{\Sigma}+RI)^{-1}\big].$ (11)
@@ 3.2 High- and Low-Temperature Behavior of the Generalization Error
$\delta(\mathbf{x})\;=\Delta_{T}(\mathbf{x})^{2}+s^{2}(\mathbf{x})+\Sigma_{l=1}^{3}(-1)^{l}\frac{C_{l}(\mathbf{x})}{t(\mathbf{x})^{l}}\prod_{i=1}^{l}\Big(1-\tfrac{i}{k}\Big)+\mathcal{O}\!\Big(t(\mathbf{x})^{-4}\Big)$ (14)
Where we have defined
@@
$\Delta_{T}(\mathbf{x}):=m(\mathbf{x})-\mu_{T}(\mathbf{x}),\quad\Delta_{R}(\mathbf{x}):=m(\mathbf{x})-\mu_{R}(\mathbf{x}),\quad t(\mathbf{x})=\frac{T}{2s(\mathbf{x})^{2}}$ (16)
and all other quantities are as in Result 1.
@@
Note that the low-temperature scaling of $\delta$ with the number of inference time samples $k$ is independent of the amount of the amount of training data. These theoretical results are compared with the experiment in Figure 6.
@@ 3.4 There exists an optimal number for inference-time samples
3.5 There exitst an optimal temperature for reward sampling
@@ 3.7 Trade-off between training and inference-time compute
Figure 6: In the plot we have chosen $S=1,\sigma=10^{-4},\gamma=10^{-3},d=10$. (a) We plot the theoretical (given in Result 3) and experimental value of $\delta$ at $T=0$ and find good agreement for $\mathbf{w}_{R}=\mathbf{w}_{T}\sim\mathcal{N}(0,2^{2}\mathbf{I})$. We see that in this domain scaling $n$ higher is less useful compared to scaling $k$. (b) We plot the theoretical (given in Result 2) and experimental value of $\delta$ and find good agreement for $\mathbf{w}_{R}=\mathbf{w}_{T}\sim\mathcal{N}(0,2^{2}\mathbf{I})$. We see that in this domain scaling $n$ higher is more useful compared to scaling $k$.
@@
$\frac{\partial\log\delta}{\partial\log k}=-2,\qquad\frac{\partial\log\delta}{\partial\log n}=-\frac{\alpha\partial_{\alpha}\big(\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}\big)}{\sigma^{2}d-2\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}},\qquad\bigg|\frac{\partial\log\delta}{\partial\log k}\bigg|\gg\bigg|\frac{\partial\log\delta}{\partial\log n}\bigg|$ (24)
@@
We empirically validate these results on advantages of inference-time scaling over training compute in Figure 6(a). It is clear from Figure 6(a) that fractional increase in $k$ decreases generalization error $\delta$ more compared to the same fractional increase in $n$ within the domain of parameters considered in the plot. However, inference-time scaling is not always advantageous over increasing training compute - we explain this in Figure 6(b). We conclude that when we have access to a good quality reward model and the task is easy enough, the addition of inference time compute is beneficial over additional training compute.
@@ 4 Qualitative agreement with large language model reasoning
In this section, we discuss implications of our theoretical results for inference time scaling of large language models.
@@
Figure 7: In the plot we have generated inference time samples from Meta-Llama-3-8B-Instruct on openai/gsm8k validation dataset (prompt included $8$ chain of thought demonstrations). For each question and response pair we used Mistral-7B-Instruct-v0.3 (with simple/weak prompt) to generate a deterministic reward score. See the associated code for the details of the prompt. For generalization error we used the definition in equation 26. The variance in the graph is due to non-zero temperature of generation in Meta-Llama-3-8B-Instruct. Plot (a), (b) shows existence of an optimal value of $k,T$ respectively.
Figure 8: We use the same setup as in Figure. 7 with two different prompts for the reward model - a simple/weak prompt and a detailed/strong prompt (see the associated code for the details of the prompt). We call these weak and strong judge respectively. Plot (a), (b) shows behavior of the optimal value of $k,T$ respectively under the change of judge. In the plot we have used a single generation from the Meta-Llama-3-8B-Instruct and kept the judge deterministic.
@@
In Figure 8 we study the change of optimal values of $k,T$ when the judge is changed from a weaker one to a stronger one by keeping the judge-LLM the same, but using a detailed prompt. During our study of the toy model, from Figure 4 we see that for a range of $c$, optimal value of $k$ remained almost fixed. However, as the reward model became stronger, i.e., $c$ is decreased, Figure 5 showed that optimal $T$ also decreased. We see similar qualitative behaviour from Figure 8(a) and 8(b) respectively.
In the linear setting, we show that the reward model that optimizes performance need not coincide with the teacher model. Leveraging properties of the trained predictor, specifically, that its mean prediction approaches the teacher’s value from below, we derive the optimal reward model. By analogy, in large language models one can employ an auxiliary classifier that learns the model’s systematic weaknesses and selects which queries should be scored by the reward model; such classifier-guided reward shaping can further improve performance. A comprehensive study of these directions is left to future work.

### v2

@@ 1 Introduction
Scaling training compute via larger models and more data drives dramatic gains (Hestness et al., 2017; Kaplan et al., 2020; Hoffmann et al., 2022). In parallel, across tasks, allowing models to ‘think longer’ at inference by sampling multiple candidates, re-ranking with a reward, or aggregating votes consistently improves precision and reliability (Wang et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Wu et al., 2024; Snell et al., 2025). Best-of-$k$ (choose the highest-reward sample based on the feedback of a judge large language model) and majority voting (choose the consensus among the generated answers) have become standard inference-time tools (Brown et al., 2024; Schaeffer et al., 2025a; Chen et al., 2024a; Huang et al., 2025a) along with self-verification (Saunders et al., 2022; Weng et al., 2023).
@@
This simple model recapitulates existing findings on large language model (LLM) inference and yields novel predictions that we evaluate in this paper. It reproduces the empirical observation that unbounded increases in the number of inference-time samples do not confer additional benefit (Snell et al., 2025). Furthermore, it predicts the existence of an optimal temperature for the reward process. Further, strong rewards shift optimal temperature to a lower value. We empirically examine and confirm this prediction using Llama-3-8B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B etc.
@@
  • In order to derive analytic insights about the inference time optimization, we derive a series expansion for $\delta$ around large $T$, making explicit its dependence on $n$, $k$, and the alignment between $\mathbf{w}_{R}$ and $\mathbf{w}_{T}$. The analytical expansion shows a sharp dependence on reward quality: when $\mathbf{w}_{R}$ is sufficiently close to $\mathbf{w}_{T}$, i.e., small $\|\mathbf{w}_{R}-\mathbf{w}_{T}\|/\|\mathbf{w}_{T}\|$, increasing $k$ monotonically decreases $\delta$, and the reward $\mathbf{w}_{R}$ that optimizes inference-time selection generally differs from the data-generating teacher $\mathbf{w}_{T}$. In contrast, when $\mathbf{w}_{R}$ is poorly aligned, $\delta$ is non-monotone in $k$, yielding an optimal finite $k$ (see Figure 1), echoing phenomena observed empirically in large language models (Snell et al., 2025). Furthermore we show that at fixed $k$, there exists an optimal temperature $T$ for the rewarding process. A similar yet distinct observation has been made in large language models (Du et al., 2025) for sampling temperature of the model itself. We emphasize that our result is about the sampling temperature of the rewarding process - different from the sampling temperature of the language model.
@@
  • To test our predictions in realistic settings, we have generated inference time samples from Llama-3-8B-Instruct and DeepSeek-R1-Distill-Qwen-1.5B on openai/gsm8k validation dataset (see Appendix G.2 for the details and more empirical results). For each question and response pair we used Mistral-7B-Instruct-v0.3 or Qwen2.5-Math-PRM-7B to generate a reward score. Finally, we observed the existence of an optimal $k,T$ similar to the theoretical predictions above (see Figure 8, 9, 10, 11 and 12). In particular for the best-of-$k$ strategy, Figure  12 shows scaling laws similar to the one observed in theoretical analysis for a nearly ideal reward model.
@@ 1.1 Related works
A growing body of work investigates how to allocate and exploit inference-time compute to improve predictive performance, with empirical gains reported across tasks and domains based on majority voting  (Chen et al., 2024a; Snell et al., 2025; Setlur et al., 2025; Arora and Zanette, 2025; Wu et al., 2024; Liu et al., 2025; Du et al., 2025; Huang et al., 2025a) or a best-of-$k$ strategy  (Wang et al., 2023; Yao et al., 2023a; Brown et al., 2024; Levi, 2024; Schaeffer et al., 2025b; Huang et al., 2025b; Chen et al., 2024b; Du et al., 2025; Chen et al., 2025). These procedures are often paired with reasoning-oriented prompting and structured search that expand the candidate set before selection (Wei et al., 2022; Yao et al., 2023b). The work of  Chen et al. (2024a) presented a theoretical model for majority voting in the premise of classification problems. More close to our work is the scaling law of Brown et al. (2024); Schaeffer et al. (2025b) for best-of-$k$ strategy. For a given trained model, both these works explain some of the empirically observed patterns at inference. We study similar questions for the regression model and discuss trade-off between training and inference time compute taking into account the quality of the reward and the sampling process.
@@ 2.2 Inference-time sampling and the reward model
$r(y,\mathbf{x})=-\left(y-\mathbf{w}_{R}\cdot\frac{\mathbf{x}}{\sqrt{d}}\right)^{2}.$ (6)
@@
$\delta=\mathbb{E}_{\mathbf{x}}(\delta(\mathbf{x})),$ (7)
$\delta(\mathbf{x})=\mathbb{E}_{y_{1},\dots,y_{k}}\left[\frac{\sum_{i=1}^{k}\big(y_{i}-\mu_{T}(\mathbf{x})\big)^{2}\,e^{-\big(y_{i}-\mu_{R}(\mathbf{x})\big)^{2}/T}}{\sum_{j=1}^{k}e^{-\big(y_{j}-\mu_{R}(\mathbf{x})\big)^{2}/T}}\right].$ (8)
@@ 3.1 Asymptotic behavior of the generalization error via deterministic equivalents
$\delta(\mathbf{x})=\mathbb{E}\left[\frac{\sum_{i=1}^{k}\big(y_{i}-\mu_{T}(\mathbf{x})\big)^{2}\,e^{-\big(y_{i}-\mu_{R}(\mathbf{x})\big)^{2}/T}}{\sum_{j=1}^{k}e^{-\big(y_{j}-\mu_{R}(\mathbf{x})\big)^{2}/T}}\right],$ (9)
to the leading order in $\sigma$. Here the expectation is over $y_{i}\sim\mathcal{N}(m(\mathbf{x}),s(\mathbf{x})^{2}),i=1,2,\dots,k$. The posterior predictive has a mean $m(\mathbf{x})$ and variance $\Sigma(\mathbf{x})=s(\mathbf{x})^{2}$ as follows
$m(\mathbf{x})=\frac{\mathbf{x}}{\sqrt{d}}^{\top}\mathbf{A}_{R}\mathbf{w}_{T},\,\,s(\mathbf{x})^{2}=\sigma^{2}+\gamma^{2}\,\frac{\mathbf{x}}{\sqrt{d}}^{\top}\mathbf{B}_{R}\,\frac{\mathbf{x}}{\sqrt{d}}.$ (10)
@@
$\mathbf{A}_{R}:=\bm{\Sigma}(\bm{\Sigma}+R\mathbf{I})^{-1},\,\,\mathbf{B}_{R}:=R(\bm{\Sigma}+R\mathbf{I})^{-1}=\mathbf{I}-\mathbf{A}_{R}.$ (11)
@@
$\hat{R}=R\Big(1-\alpha\,m_{\bm{\Sigma}}(R)\Big)=\frac{\sigma^{2}}{\gamma^{2}}\alpha,\,\,m_{\bm{\Sigma}}(R):=\frac{1}{d}\text{Tr}\!\big[\bm{\Sigma}(\bm{\Sigma}+RI)^{-1}\big].$ (12)
@@ 3.2 High- and low-temperature behavior of the generalization error
$\delta(\mathbf{x})$ $=\Delta_{T}(\mathbf{x})^{2}+s^{2}(\mathbf{x})$ (15)
$+\Sigma_{l=1}^{3}(-1)^{l}\frac{C_{l}(\mathbf{x})}{t(\mathbf{x})^{l}}\prod_{i=1}^{l}\Big(1-\tfrac{i}{k}\Big)+\mathcal{O}\!\Big(t(\mathbf{x})^{-4}\Big)$
Where we have defined $t(\mathbf{x})=\frac{T}{2s(\mathbf{x})^{2}}$ and
@@
$\Delta_{T}(\mathbf{x})=m(\mathbf{x})-\mu_{T}(\mathbf{x}),\quad\Delta_{R}(\mathbf{x})=m(\mathbf{x})-\mu_{R}(\mathbf{x}).$
All other quantities are as in Result 1.
@@
Note that the low-temperature scaling of $\delta$ with the number of inference time samples $k$ is independent of the amount of the amount of training data. These theoretical results are compared with the experiment in Figure 6, Figure 7.
@@ 3.4 There exists an optimal number for inference-time samples
3.5 There exists an optimal temperature for reward sampling
@@ 3.7 Trade-off between training and inference-time compute
Figure 6: In the plot we have chosen $S=1,\sigma=10^{-4},\gamma=10^{-3},d=10$. We plot the theoretical (given in Result 2) and experimental value of $\delta$ and find good agreement for $\mathbf{w}_{R}=\mathbf{w}_{T}\sim\mathcal{N}(0,2^{2}\mathbf{I})$. We see that in this domain scaling $n$ higher is more useful compared to scaling $k$.
Figure 7: In the plot we have chosen $S=1,\sigma=10^{-4},\gamma=10^{-3},d=10$. We plot the theoretical (given in Result 3) and experimental value of $\delta$ at $T=0$ and find good agreement for $\mathbf{w}_{R}=\mathbf{w}_{T}\sim\mathcal{N}(0,2^{2}\mathbf{I})$. We see that in this domain scaling $n$ higher is less useful compared to scaling $k$.
@@
$\frac{\partial\log\delta}{\partial\log k}=-2,\qquad\frac{\partial\log\delta}{\partial\log n}=-\frac{\alpha\partial_{\alpha}\big(\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}\big)}{\sigma^{2}d-2\mathbf{u}^{\top}\bm{\Sigma}\mathbf{u}}$ (24)
and hence
$\bigg|\frac{\partial\log\delta}{\partial\log k}\bigg|\gg\bigg|\frac{\partial\log\delta}{\partial\log n}\bigg|$ (25)
@@
We empirically validate these results on advantages of inference-time scaling over training compute in Figure 7. It is clear from Figure 7 that fractional increase in $k$ decreases generalization error $\delta$ more compared to the same fractional increase in $n$ within the domain of parameters considered in the plot. However, inference-time scaling is not always advantageous over increasing training compute - we explain this in Figure 6. We conclude that when we have access to a good quality reward model and the task is easy enough, the addition of inference time compute is beneficial over additional training compute.
@@ 4 Qualitative agreement with large language model reasoning
Figure 8: In the plot we have generated inference time samples from Llama-3-8B-Instruct. For each question and response pair we used Mistral-7B-Instruct-v0.3 (with simple/weak prompt). For generalization error we used the definition in equation 27. The variance in the graph is due to non-zero temperature of generation in Llama-3-8B-Instruct. Plot shows existence of an optimal value of $k$.
In this section, we discuss implications of our theoretical results for inference time scaling of large language models such as Llama-3-8B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B on openai/gsm8k validation dataset (prompt included $8$ chain of thought demonstrations). As the Judge model we used Mistral-7B-Instruct-v0.3 or Qwen2.5-Math-PRM-7B. For the details of experimental design see Appendix G.2.
Figure 9: We use the same setup as in Figure. 8. Plot shows existence of an optimal value of $T$.
@@
In Figure 10 and Figure 11 we study the change of optimal values of $k,T$ when the judge is changed from a weaker one to a stronger one by keeping the judge-LLM the same, but using a detailed prompt. During our study of the toy model, from Figure 4 we see that for a range of $c$, optimal value of $k$ remained almost fixed. However, as the reward model became stronger, i.e., $c$ is decreased, Figure 5 showed that optimal $T$ also decreased. We see similar qualitative behavior from Figure 10 and 11 respectively.
Figure 10: We use the same setup as in Figure. 8 with two different prompts for the reward model - a simple/weak prompt and a detailed/strong prompt (see Appendix G.2 for the details of the prompt). We call these weak and strong judge respectively. Plot shows behavior of the optimal value of $k$ respectively under the change of judge. In the plot we have used a single generation from the Llama-3-8B-Instruct and kept the judge deterministic.
Figure 11: We use the same setup as in Figure 8 with two different prompts for the reward model as in Figure 10. Plot shows behavior of the optimal value of $T$ under the change of judge.
Figure 12: In this plot we plot the error $1+\delta$ of various base models vs $k$ at fixed reward temperature $T=0$ for nearly ideal reward model Qwen2.5-Math-PRM-7B. It is clear from the plot that the curves are well fitted by an effective scaling law similar to the result in Remark 5 except that the exponent of $k$ is different.
In Figure 12 we use Qwen2.5-Math-PRM-7B as the reward model. This verifies each step of the reasoning and corresponds to nearly ideal reward. The plot shows that for best-of-$k$ method the error decreases as a power law in inference time samples (almost monotonically). Similar result is established theoretically in Remark 5. Note that the judge in our case takes into account the reasoning process, not just final result of being correct or wrong and our metric is much different compared to pass@k used in Brown et al. (2024); Schaeffer et al. (2025a).
5 Conclusion
We introduce a simple, solvable model to study how inference-time scaling works, where multiple candidate answers are generated and chosen using a reward-based selection process. Our analysis shows that when the reward model is well aligned with the task, increasing the number of inference time samples reliably improves generalization, but when it is misaligned, there is an optimal finite number of samples and an optimal temperature for selection. In the best-of-k setting with an ideal reward model, we identify a power law of generalization error and demonstrate when extra inference-time compute is more valuable than additional training. Finally, our experiments with modern language models on a math problem benchmark validate theoretical results.
In the linear setting, we show that the reward model that optimizes performance need not coincide with the teacher model. Leveraging properties of the trained predictor, we derive the optimal reward model. By analogy, in large language models one can employ an auxiliary classifier that learns the model’s systematic weaknesses and selects which queries should be scored by the reward model; such classifier-guided reward shaping can further improve performance. A comprehensive study of these directions is left to future work.